A number is 5 times another number. The product of the two numbers are 245. What are the two numbers?
Answer by tvandenberg(45)

Y = 5X
X*Y = 245
X * (5X) = 245
X^2 = 245 / 5 = 49
X = SQRT(49) = 7
Y = 5*7 = 35